# Planar Collisionless Shock Simulations with Semi-Implicit Particle-in-Cell Model FLEKS

### v1

@@ Abstract
This study investigates the applicability of the semi-implicit particle-in-cell code FLEKS to collisionless shock simulations, with a focus on the parameter regime relevant to global magnetosphere modeling. We examine one- and two-dimensional local planar shock simulations, initialized using MHD states with upstream conditions representative of the solar wind at 1 au, for both quasi-perpendicular and quasi-parallel configurations. The refined algorithm in FLEKS proves robust, enabling accurate shock simulations with a grid resolution on the order of the electron inertial length $d_{e}$. Our simulations successfully capture key shock features, including shock structures (foot, ramp, overshoot, and undershoot), upstream and downstream waves (fast magnetosonic, whistler, Alfvén ion-cyclotron, and mirror modes), and non-Maxwellian particle distributions. Crucially, we find that at least two spatial dimensions are critical for accurately reproducing downstream wave physics in quasi-perpendicular shocks and capturing the complex dynamics of quasi-parallel shocks, including surface rippling, shocklets, SLAMS, and jets. Furthermore, our parameter studies demonstrate the impact of mass ratio and grid resolution on shock physics. This work provides valuable guidance for selecting appropriate physical and numerical parameters for a semi-implicit PIC code, paving the way for incorporating kinetic shock processes into large-scale space plasma simulations with the MHD-AEPIC model.
@@ 1 Introduction
Collisionless shocks are ubiquitous in space plasmas, marking abrupt transitions in the velocity of charged particles. Unlike shocks in neutral fluids, collisionless shocks are mediated by electromagnetic fields rather than particle collisions. These shocks play a crucial role in diverse astrophysical environments, including astrophysical jets, the solar wind, the heliosphere, and planetary magnetospheres (see A. Balogh & R. A. Treumann [2013] and references therein). The passage of a shock can dramatically alter the properties of a plasma, leading to particle acceleration, energy dissipation, and the generation of various waves.
@@
Recent developments in PIC methods have led to semi-implicit algorithms that overcome some of these limitations. These methods relax the stringent stability constraints of explicit PIC while avoiding the nonlinear iterations required in fully implicit PIC [G. Lapenta, 2023; J. Ren & G. Lapenta, 2024]. Semi-implicit PIC methods offer improved energy conservation and numerical stability, enabling simulations with larger grid cells and time steps. This study investigates the applicability of a semi-implicit, energy and charge conserving PIC model FLEKS [Y. Chen et al., 2023] to collisionless shock simulations, using parameters relevant to global terrestrial magnetosphere modeling. FLEKS offers several advantages, including near-exact energy and charge conservation [G. Lapenta, 2017; Y. Chen & G. Tóth, 2019] and the ability to use larger grid sizes and time steps compared to explicit PIC methods. Furthermore, we have proposed to address the numerical instabilities encountered in the shock simulations by (a) adding a Lax-Friedrichs-type diffusion term to the electromagnetic field solvers to suppress spurious oscillations near discontinuities and (b) using a novel particle current calculation method to reduce statistical noise (Chen et al., 2025, in prep.).
@@
In this study, we present one-dimensional (1D) and two-dimensional (2D) local planar shock simulations using FLEKS within the Space Weather Modeling Framework (SWMF, G. Tóth et al. [2012]). We explore both quasi-perpendicular and quasi-parallel shock configurations, varying the ion-to-electron mass ratio $m_{i}/m_{e}$ and grid resolution $\Delta x$ on the order of electron skin depth $d_{e}$. Our results demonstrate that the refined FLEKS algorithm accurately reproduces key shock features, including shock structures, wave generation (fast magnetosonic, whistler, Alfvén ion-cyclotron, and mirror modes), and non-Maxwellian particle distributions, provided that the resolution is on a sub-ion scale and a 2D spatial domain is used. These findings validate FLEKS for collisionless shock simulations and offer guidance for model parameter selection. Ultimately, this work paves the way for incorporating kinetic shock physics into large-scale magnetosphere models.
@@ 2 Methodology
The FLexible Exascale Kinetic Simulator (FLEKS) is a semi-implicit PIC code that utilizes a Gauss’s Law-satisfying Energy-Conserving Semi-Implicit Particle-in-Cell method (GL-ECSIM). In FLEKS, Maxwell’s equations are solved on a collocated staggered grid, where electric fields are at grid nodes and magnetic fields at cell centers. To address spurious oscillations near discontinuities and statistical noise in fast plasma flows, FLEKS incorporates two key numerical techniques. First, a Lax-Friedrichs-type diffusion term with a flux limiter is introduced into the Maxwell solver to suppress unphysical oscillations. Second, a novel approach calculates the current density in the comoving frame, which significantly reduces particle noise in simulations with fast plasma flows. Digital filter smoothing is also applied to the comoving frame current density to further enhance numerical stability and mitigate numerical cooling effects. Readers interested in more details are referred to Chen et al. (2025, in preparation).
Collisionless shocks in space are characterized by several dimensionless parameters, including the Alfvén Mach number ($\text{M}_{\text{A}}$) and the electron and ion plasma betas ($\beta_{e}$ and $\beta_{i}$). These parameters govern the fundamental properties of shocks and can vary over orders of magnitude. This study investigates upstream plasma conditions typical of the solar wind at 1 au near Earth’s magnetosphere. We selected an average plasma state with $\text{M}_{\text{A}}=7$, $\beta_{i}=0.5$, and $\beta_{e}=1.0$, based on solar wind temperature statistics from NASA’s Wind spacecraft as reported by C. S. Salem et al. [2023] and summarized in their Figure 9 for the dependence on solar wind speed and $\beta$.
To reduce computational cost, we employed an artificially reduced proton-to-electron mass ratio of $m_{i}/m_{e}=25$ and a ratio of electron plasma frequency to electron cyclotron frequency of $\omega_{pe}^{2}/\omega_{ce}^{2}=10$. Since $\omega_{pe}/\omega_{ce}=c/V_{Ae}$, where $V_{Ae}=B/\sqrt{\mu_{0}\rho_{e}}=V_{Ai}\sqrt{m_{i}/m_{e}}$ is the electron Alfvén speed, these ratios determine the effective speed of light and other key plasma scales in the simulation. With an upstream plasma number density of $n=n_{i}=n_{e}=3\,\text{cm}^{-3}$ and magnetic field of $B=5\,\text{nT}$, the system is fully determined. Table 1 compares the electron and ion spatiotemporal scales in the simulation to their corresponding values in nature.
[…]
@@
Table 2: List of planar shock runs. $\theta_{Bn,\text{up}}$ is the upstream shock-normal angle and $\Delta x$ is the grid resolution normalized by the upstream electron inertial length.
Run Dimension $\theta_{Bn,\text{up}}\,[^{\circ}]$ $m_{i}/m_{e}$ $\Delta x\,[d_{e,\text{up}}]$
1 1 87 25 1
2 1 30 25 1
3 1 87 100 1
4 1 87 400 1
5 1 30 100 1
6 1 30 400 1
7 1 87 25 0.5
8 1 87 25 2
9 1 87 25 4
10 1 30 25 0.5
11 1 30 25 2
12 1 30 25 4
13 2 87 25 1
14 2 30 25 1
@@
Unlike global magnetosphere simulations, which often use the planetary radius as a characteristic length scale, local shock simulations lack an inherent spatial reference. Therefore, we typically normalize quantities to the relevant kinetic scales of ions and electrons. To emphasize the scaling relations in the local system, all displayed plasma quantities are normalized to characteristic upstream ion scales: $d_{i,\text{up}}$ for length, $\omega_{ci,\text{up}}$ for frequency, $n_{i,\text{up}}$ for density, $V_{Ai,\text{up}}$ for velocity, and $P_{i,\text{up}}$ for pressure, unless otherwise specified. Electromagnetic fields are normalized to the magnitudes of the upstream values $B_{\text{up}}$ and $E_{\text{up}}=|\mathbf{V}_{Ai,\text{up}}\times\mathbf{B}_{\text{up}}|$. All the displayed times are normalized by the upstream proton gyrofrequency $\omega_{ci}$, where a $2\pi$ factor is ignored for converting from gyrofrequencies to periods. The output cadence is set to $1\,\text{s}\simeq 0.08\,\omega_{ci}^{-1}$.
@@
As a PIC model, FLEKS unavoidably exhibits statistical noise due to the use of a finite number of macro-particles that scales as $\propto N^{-1/2}$, where $N$ is the number of macro-particles per cell. Empirically we find a noise level $\sim 10\%$ for $N=100$ and $\sim 5\%$ for $N=400$ from a 1D free-stream test with uniform plasma states, electromagnetic fields, and periodic boundary conditions. In the 1D and 2D shock simulations presented here, we used $N=400$, a value that remains feasible for global 2D magnetosphere simulations while effectively suppressing noise. While techniques like high-order shape functions [S. Muralikrishnan et al., 2021] or kernel density estimation [W. Wu & H. Qin, 2018] can help mitigate noise, they introduce additional complexities in performance and implementation. Therefore, we currently employ the standard second-order scheme with a triangular shape function. To avoid visual bias in velocity space plots, we applied an ion cutoff density value of $w_{\text{min}}=10^{-3}w_{\text{peak}}$ for quasi-perpendicular shock simulations and $w_{\text{min}}=10^{-4}w_{\text{peak}}$ for quasi-parallel shock simulations, where $w_{\text{peak}}$ is the maximum phase density in the sampled region. Phase space densities below these thresholds were set to 0 and displayed as white in the logarithmic scale colormaps. A 3-point sliding window smoothing is applied for showing the plasma moments in the line profiles.
@@ 2.1 Local 1D Simulations
Figure 1: 1D simulation of a quasi-perpendicular shock with $\theta_{Bn}=87^{\circ}$ and $m_{i}/m_{e}=25$. The plasma moments and EM fields are normalized by the upstream quantities. Phase space distributions are normalized by the maximum phase space density in the displayed regions. (a-e) Close-up of plasma quantities and electromagnetic fields near the shock front at $t=20\,\omega_{ci}^{-1}$. (f) $x-t$ stack plot of the total magnetic field, overlaid with ion density contours. (g,h) Ion and electron $x-v_{x}$ phase space plot, with the $B_{z}$ profile shown in blue for reference. Ion and electron velocity space distributions are displayed in the downstream region (i-l), downstream right behind the shock front (m-p), and upstream right next to the shock front (q-t).
Figure 1 presents a detailed view of plasma moments, electromagnetic fields, and electron and ion phase space distributions near the quasi-perpendicular $\theta_{Bn}=87^{\circ}$ shock at $t=20\,\omega_{ci}^{-1}$, with the shock front located around $x=5\,d_{i,\text{up}}$. The characteristic shock structures, including the foot, ramp, overshoot, and undershoot, are clearly discernible in the density (Figure 1a) and magnetic field (Figure 1d) profiles. The ramp width, defined as the transition from the onset of magnetic field variation to the first maximum (overshoot), measures approximately 0.5 ion convective gyroradii $r_{ci,\text{up}}=V_{\text{up}}/\omega_{ci,\text{up}}$, or 4 upstream ion inertial lengths $d_{i,\text{up}}=c/\omega_{pi,\text{up}}$. This value aligns with observations from early AMPTE satellite measurements [S. N. Walker et al., 1999a, b] and subsequent Cluster and THEMIS data (Figure 2 in Y. OHobara et al. [2010]) for narrow ramps at $\text{M}_{\text{A}}=7$. Due to shock drift acceleration by the upstream convection electric field (primarily $E_{y}$ in Figure 1e), reflected protons exhibit higher energy than directly transmitted protons. Their gyration downstream results in arches and holes in the $x-V_{ix}$ phase space (Figure 1g) at large $V_{ix}$ magnitudes, although these features are partially smoothed by the thermal velocity. The cross-shock electric potential, arising from charge separation due to the abrupt magnetic field jump, can be estimated from the electric field $E_{x}$ in Figure 1e as $\Delta\phi=\int E_{x}\mathrm{d}x\simeq 430\,\text{V}$. This yields an electric potential energy $U_{E}=q_{i}\Delta\phi\approx 0.4E_{k}$, where $E_{k}=m_{i}V_{i,\text{up}}^{2}/2$ is the upstream ion kinetic energy. Consistent with the ion reflection model proposed in the Appendix of Y. V. Khotyaintsev et al. [2024], this scenario falls into the category where ions, upon reflection, are decelerated by $\Delta\phi$ and then undergo cyclotron gyration due to the increased magnetic field in the ramp, leading to a broad distribution of the reflected population within the ramp.
@@
[…]
@@
Correlated periodic downstream fluctuations are observed in the plasma density, total magnetic field, and electron thermal pressures (Figure 1a, 1c, and 1d). These postshock oscillations are not indicative of transmitted or generated waves. Instead, they are periodic structures that arise from ion ring distributions. This is true despite their exhibiting density and magnetic field correlations similar to compressional fast magnetosonic waves [C. T. Russell et al., 2009]. Interpretations of these structures as damping solitons [R. Sagdeev, 1966; C. Kennel, 1967] are inconsistent with the shock’s $\beta$ and Mach number. These structures also differ from the stationary downstream oscillations analyzed in L. Ofman et al. [2009]; L. Ofman & M. Gedalin [2013], where subcritical shocks maintain total pressure balance with anti-correlated magnetic field and ion pressure variations. While their descriptions involve non-gyrotropic downstream ions, our simulations clearly show a symmetric velocity distribution in the perpendicular $V_{ix}-V_{iy}$ plane. Ion gyration and the formation of downstream ring distributions (Figure 1i and 1m) are intrinsic consequences of ion dynamics across the shock front in kinetic models, analogous to overshoot formation in supercritical shocks. In the downstream region, most ions directly penetrate the shock front and are decelerated by the cross-shock potential. The fraction of reflected ions is small (Figure 1g), limited to a normal length of $2\,d_{i,\text{up}}$ in the ramp, consistent with low to moderate Mach number ($\text{M}_{\text{A}}\leq 10$) quasi-perpendicular shocks [A. Balogh & R. A. Treumann, 2013; D. B. Graham & Y. V. Khotyaintsev, 2025]. Minimal ion deflection by the magnetic field occurs due to the ramp width being much smaller than the upstream ion convective gyroradius $r_{ci,\text{up}}=V_{i,\text{up}}/\omega_{ci,\text{up}}$. These non-isotropic reflected ions (Figure 1g, 1q-r), after being drift-accelerated along the y-direction, recross the ramp and form a downstream gyrating beam [M. M. Leroy et al., 1982; D. Burgess et al., 1989], leading to oscillating ion pressure. The spatial scale of the periodic structure $L$ corresponds to the ion ring distribution in Figure 1i, centered at $V_{ix}=-2V_{Ai,\text{up}},V_{iy}=0$ with a radius of $V_{ring}\simeq 8V_{Ai,\text{up}}$, where $L\sim 2.4\,d_{i,\text{up}}\sim m_{i}V_{ring}/(q_{i}B_{\text{down}})$). Due to their frozen-in behavior with the background magnetic field, the total magnetic field also exhibits coherent periodicity (Figure 1d). With suppressed wave-particle interactions due to the 1D configuration and slow gyrophase mixing, this postshock periodic structure persists nearly to the downstream simulation boundary.
@@
We conducted a companion simulation with $\Delta x=0.5\,d_{e}$ and confirmed that the downstream periodic structures are ion-scale phenomena, independent of electron-scale numerical convergence. However, the inclusion of electron kinetics in FLEKS introduces electrostatic electron pressure modulation by ion and magnetic pressure variations, evident in the correlated spatial oscillation of the $V_{ex}$ extent (Figure 1h) and the modulated pressure components $P_{exx}$, $P_{eyy}$, and $P_{ezz}$ (Figure 1c and 1h). This effect is absent in hybrid models that treat electrons as a massless fluid in an ideal thermodynamic process (e.g. adiabatic if $\gamma=5/3$; isothermal if $\gamma=1$).

### v2

@@ Abstract
This study investigates the applicability of the semi-implicit particle-in-cell code FLEKS to heliospheric shock simulations. We examine one- and two-dimensional local planar shock simulations, initialized using MHD states with upstream conditions representative of plasmas in the hypersonic, $\beta\sim 1$ regime, for both quasi-perpendicular and quasi-parallel configurations. The refined algorithm in FLEKS proves robust, enabling accurate shock simulations with a grid resolution on the order of the electron inertial length $d_{e}$. Our simulations successfully capture key shock features, including shock structures (foot, ramp, overshoot, and undershoot), upstream and downstream waves (fast magnetosonic, whistler, Alfvén ion-cyclotron, and mirror modes), and non-Maxwellian particle distributions. Crucially, we find that at least two spatial dimensions are critical for accurately reproducing downstream wave physics in quasi-perpendicular shocks and capturing the complex dynamics of quasi-parallel shocks, including surface rippling, shocklets, SLAMS, magnetic reconnection and jets. Furthermore, our parameter studies demonstrate the impact of mass ratio and grid resolution on shock physics. This work provides valuable guidance for selecting appropriate physical and numerical parameters for shock simulations using a semi-implicit PIC method, paving the way for incorporating kinetic shock processes into large-scale collisionless plasma simulations with the MHD-AEPIC model.
@@ 1 Introduction
Collisionless shocks are ubiquitous in space plasmas, marking abrupt transitions in the velocity of charged particles. Unlike shocks in neutral fluids, collisionless shocks are mediated by electromagnetic fields rather than particle collisions. These shocks play a crucial role in diverse astrophysical environments, including astrophysical jets [M. Livio, 1999; R. Blandford et al., 2019], the solar wind [D. M. Oliveira, 2023], the heliosphere [R. G. Stone & B. T. Tsurutani, 1985], and planetary magnetospheres (see more in A. Balogh & R. A. Treumann [2013] and references therein). The passage of a shock can dramatically alter the properties of a plasma, leading to particle acceleration, energy dissipation, and the generation of various wave modes across a broad spectrum of frequencies.
@@
Recent developments in PIC methods have led to semi-implicit algorithms that offer a promising pathway to overcome some of these long-standing limitations. These methods relax the stringent stability constraints of explicit PIC while avoiding the computationally expensive nonlinear iterations required in fully implicit PIC [G. Lapenta, 2023; J. Ren & G. Lapenta, 2024]. Semi-implicit PIC methods, by allowing for larger grid cells and time steps, offer improved energy conservation and numerical stability, making them more amenable to simulating the larger physical domains characteristic of astrophysical scenarios. This study investigates the applicability of a semi-implicit, energy and charge conserving PIC model, FLEKS [Y. Chen et al., 2023], to collisionless shock simulations. FLEKS offers several key advantages, including near-exact energy and charge conservation [G. Lapenta, 2017; Y. Chen & G. Tóth, 2019] and the ability to use significantly larger grid sizes and time steps compared to traditional explicit PIC methods. However, for shock tests with the original semi-implicit algorithm, we found that the super-Alfvénic upstream region shows unstable wave growth non-existing in nature. These so-called ”spurious” oscillations are essentially associated with particle noise that gets amplified in the solution of the electric field from the semi-implicit solver.
@@
We have proposed to address numerical instabilities encountered in shock simulations by (a) adding a Lax-Friedrichs-type diffusion term to the electromagnetic field solvers to suppress spurious oscillations near discontinuities and (b) using a novel particle current calculation method to reduce statistical noise [Y. Chen et al., 2025]. In this study, we present one-dimensional (1D) and two-dimensional (2D) local planar shock simulations using the refined FLEKS algorithm within the Space Weather Modeling Framework (SWMF, G. Tóth et al. [2012]). The simulations focus on non-relativistic, heliospheric shocks, typically characterized by Alfvén Mach numbers $\mathrm{M}_{\mathrm{A}}\lesssim\mathcal{O}(10)$. Astrophysical shocks, such as those associated with supernova remnants, can exhibit much higher Mach numbers, often reaching tens or hundreds. In such high-$\mathrm{M}_{\mathrm{A}}$ regimes, different kinetic instabilities, like the Buneman instability [Y. Matsumoto et al., 2012], Weibel instability [J. Kropotina et al., 2023], and Bell instability [J. Park et al., 2015], may become dominant in structuring the shock front and adjacent regions. This sensitivity of plasma physics to the specific parameter regime underscores the importance of conducting targeted simulations to investigate phenomena within distinct environments. For heliospheric shocks, we explore both quasi-perpendicular and quasi-parallel configurations, varying the ion-to-electron mass ratio $m_{i}/m_{e}$ and grid resolution $\Delta x$ on the order of the electron skin depth $d_{e}$. Our results demonstrate that the refined FLEKS algorithm accurately reproduces key hypersonic shock features with conditions representative at 1 AU, including the detailed shock structures, the generation of various wave modes (e.g., fast magnetosonic, whistler, Alfvén ion-cyclotron, and mirror modes), and the formation of non-Maxwellian particle distributions, provided that the resolution is on a sub-ion scale and a 2D spatial domain is used. These findings validate the enhanced FLEKS capabilities for collisionless shock simulations and offer crucial guidance for model parameter selection in future studies. Ultimately, this work represents a significant step towards incorporating detailed kinetic shock physics into large-scale, global plasma models, with direct implications for understanding particle acceleration and energy conversion in diverse astrophysical environments beyond just terrestrial applications.
@@ 2 Methodology
The FLexible Exascale Kinetic Simulator (FLEKS) is a semi-implicit PIC code that utilizes a Gauss’s Law-satisfying Energy-Conserving Semi-Implicit Method (GL-ECSIM). In FLEKS, Maxwell’s equations are solved on a collocated staggered grid, where electric fields are at grid nodes and magnetic fields at cell centers. To address spurious oscillations near discontinuities and statistical noise in fast plasma flows, recent FLEKS version incorporates two key numerical techniques. First, a Lax-Friedrichs-type diffusion term with a flux limiter is introduced into the Maxwell solver to suppress unphysical oscillations. Second, a novel approach calculates the current density in the comoving frame, which significantly reduces particle noise in simulations with fast plasma flows. Digital filter smoothing is also applied to the comoving frame current density to further enhance numerical stability and mitigate numerical cooling effects. Readers interested in more details are referred to Y. Chen et al. [2025].
Collisionless shocks in space are characterized by several dimensionless parameters, including the Alfvén Mach number ($\text{M}_{\text{A}}$) and the electron and ion plasma betas ($\beta_{e}$ and $\beta_{i}$). These parameters govern the fundamental properties of shocks and can vary over orders of magnitude across all kinds of astrophysical shocks. This study primarily investigates upstream plasma conditions typical of the solar wind at 1 AU near the terrestrial magnetosphere. We selected an average plasma state with $\text{M}_{\text{A}}=7$, $\beta_{i}=0.5$, and $\beta_{e}=1.0$, based on solar wind temperature statistics from NASA’s Wind spacecraft as reported by C. S. Salem et al. [2023] and summarized in their Figure 9 for the dependence on solar wind speed and $\beta$. Note that similar parameters in the hypersonic plasma regime can occur at other solar system planets (e.g. Mars [D. Liu et al., 2021]) and also astrophysical systems, which presents a broader scope outside Earth.
To reduce computational cost, we employed an artificially reduced proton-to-electron mass ratio of $m_{i}/m_{e}=25$ and a ratio of electron plasma frequency to electron cyclotron frequency of $\omega_{pe}^{2}/\omega_{ce}^{2}=10$. Since $\omega_{pe}/\omega_{ce}=c/V_{Ae}$, where $V_{Ae}=B/\sqrt{\mu_{0}\rho_{e}}=V_{Ai}\sqrt{m_{i}/m_{e}}$ is the electron Alfvén speed, these ratios determine the effective speed of light and other key plasma scales in the simulation. With an upstream plasma number density of $n=n_{i}=n_{e}=3\,\text{cm}^{-3}$ and magnetic field of $B=5\,\text{nT}$, the system is fully determined. Table 1 compares the electron and ion spatiotemporal scales in the simulation to their corresponding values in nature. Here, $\lambda_{D}$ is the Debye length, $d_{i}$ is the ion inertial length, $r_{ci}$ is the ion gyroradius, $d_{e}$ is the electron inertial length, $r_{ce}$ is the electron gyroradius, $\omega_{pe}$ is the electron gyrofrequency, $\omega_{ci}$ is the ion cyclotron frequency, and $\omega_{pi}$ is the ion gyrofrequency. In the following context, all the characteristic quantities are referred to the upstream values (”up”) unless otherwise specified.
Table 1: Plasma characteristic scales upstream of the 1D and 2D simulations assuming $m_{i}/m_{e}=25$, $\beta_{i}=0.5$, $\beta_{e}=1$, $n_{i}=n_{e}=3\,\text{cm}^{-3}$, $B=5\,\text{nT}$. The simulated speed of light is set to $c=3\times 10^{6}\,\text{m/s}$, which is 1/100 the vacuum speed of light.
@@
Table 2: List of planar shock runs. $\theta_{Bn,\text{up}}$ is the upstream shock-normal angle, $\mathrm{M}_{\mathrm{A,up}}$ is the upstream Mach number, $m_{i}/m_{e}$ is the mass ratio, and $\Delta x$ is the grid resolution normalized by the upstream electron inertial length.
Run Dimension $\theta_{Bn,\text{up}}\,[^{\circ}]$ $\mathrm{M}_{\mathrm{A,up}}$ $m_{i}/m_{e}$ $\Delta x\,[d_{e,\text{up}}]$
1 1 87 7 25 1
2 1 30 7 25 1
3 1 87 7 100 1
4 1 87 7 400 1
5 1 30 7 100 1
6 1 30 7 400 1
7 1 87 7 25 0.5
8 1 87 7 25 2
9 1 87 7 25 4
10 1 30 7 25 0.5
11 1 30 7 25 2
12 1 30 7 25 4
13 2 87 7 25 1
14 2 30 7 25 1
15 2 30 12 25 1
16 2 0 7 25 1
@@
Unlike global magnetosphere simulations, which often use the planetary radius as a characteristic length scale, local shock simulations lack an inherent spatial reference. Therefore, we typically normalize quantities to the relevant kinetic scales of ions and electrons. To emphasize the scaling relations in the local system, all displayed plasma quantities are normalized to characteristic upstream ion scales: $d_{i,\text{up}}$ for length, $\omega_{ci,\text{up}}$ for frequency, $n_{i,\text{up}}$ for density, $V_{Ai,\text{up}}$ for velocity, and $P_{i,\text{up}}$ for pressure, unless otherwise specified. Electromagnetic fields are normalized to the magnitudes of the upstream values $B_{\text{up}}$ and $E_{\text{up}}=|\mathbf{V}_{Ai,\text{up}}\times\mathbf{B}_{\text{up}}|$. All the displayed times are normalized by the upstream proton gyrofrequency $\omega_{ci,\text{up}}$, where a $2\pi$ factor is ignored for converting from gyrofrequencies to periods. The output cadence is set to $1\,\text{s}\simeq 0.08\,\omega_{ci,\text{up}}^{-1}$.
@@
As a PIC model, FLEKS unavoidably exhibits statistical noise due to the use of a finite number of macro-particles that scales as $\propto N^{-1/2}$, where $N$ is the number of macro-particles per cell. Empirically we find a noise level $\sim 10\%$ for $N=100$ and $\sim 5\%$ for $N=400$ from a 1D free-stream test with uniform plasma states, electromagnetic fields, and periodic boundary conditions. In the 1D and 2D shock simulations presented here, we used $N=400$, a value that remains feasible for global 2D magnetosphere simulations while effectively suppressing noise. While techniques like high-order shape functions [S. Muralikrishnan et al., 2021] or kernel density estimation [W. Wu & H. Qin, 2018] can help mitigate noise, they introduce additional complexities in performance and implementation. Therefore, we currently employ the standard second-order scheme with a triangular shape function. To avoid visual bias in velocity space plots, we applied an ion cutoff density value of $w_{\text{min}}=10^{-3}w_{\text{peak}}$ for quasi-perpendicular shock simulations and $w_{\text{min}}=10^{-4}w_{\text{peak}}$ for quasi-parallel shock simulations, where $w_{\text{peak}}$ is the maximum phase space density in the sampled region. Phase space densities below these thresholds were set to 0 and displayed as white in the logarithmic scale colormaps. A 3-point sliding window smoothing is applied for showing the plasma moments in the line profiles.
3 Results
@@ 3.1 Local 1D Simulations
Figure 1: 1D simulation of a quasi-perpendicular shock with $\theta_{Bn}=87^{\circ}$ and $m_{i}/m_{e}=25$. The plasma moments and EM fields are normalized by the upstream quantities. Phase space distributions are normalized by the maximum phase space density in the displayed regions. (a-e) Close-up of plasma quantities and electromagnetic fields near the shock front at $t=20\,\omega_{ci,\text{up}}$. (f) xtx-t stack plot of the total magnetic field, overlaid with ion density contours. (g,h) Ion and electron $x-v_{x}$ phase space plot, with the $B_{z}$ profile shown in blue for reference. Ion and electron velocity space distributions are displayed in the downstream region (i-l), downstream right behind the shock front (m-p), and upstream right next to the shock front (q-t).
Figure 1 presents a detailed view of plasma moments, electromagnetic fields, and electron and ion phase space distributions near the quasi-perpendicular $\theta_{Bn}=87^{\circ}$ shock at $t=20\,\omega_{ci,\text{up}}$, with the shock front located around $x=5\,d_{i,\text{up}}$. The characteristic shock structures, including the foot, ramp, overshoot, and undershoot, are clearly discernible in the density (Figure 1a) and magnetic field (Figure 1d) profiles. The ramp width, defined as the transition from the onset of magnetic field variation to the first maximum (overshoot), measures approximately 0.5 ion convective gyroradii $r_{ci,\text{up}}=V_{\text{up}}/\omega_{ci,\text{up}}$, or 4 upstream ion inertial lengths $d_{i,\text{up}}=c/\omega_{pi,\text{up}}$.  Due to shock drift acceleration by the upstream convection electric field (primarily $E_{y}$ in Figure 1e), reflected protons exhibit higher energy than directly transmitted protons. Their gyration downstream results in arches and holes in the $x-V_{ix}$ phase space (Figure 1g) at large $V_{ix}$ magnitudes, although these features are partially smoothed by the thermal velocity. The cross-shock electric potential, arising from charge separation due to the abrupt magnetic field jump, can be estimated from the electric field $E_{x}$ in Figure 1e as $\Delta\phi=\int E_{x}\mathrm{d}x\simeq 430\,\text{V}$. This yields an electric potential energy $U_{E}=q_{i}\Delta\phi\approx 0.4E_{k}$, where $E_{k}=m_{i}V_{i,\text{up}}^{2}/2$ is the upstream ion kinetic energy.
@@
Although the initial MHD conditions for our PIC simulations satisfy the stationarity requirement, i.e., pressure balance between magnetic, ram, and thermal pressures ($P_{B}+P_{\text{ram}}+P_{\text{th}}=\text{const.}$) across the shock front, this balance is disrupted as the shock evolves under the fully kinetic solver in the supercritical regime ($M_{A}>M_{c}\sim 2.5$, where $M_{c}$ is the first critical Mach number described in M. M. Leroy et al. [1982]). In this simulation, the shock front moves less than $10\,d_{i,\text{up}}\simeq 1320\,\text{km}$ in $20\,\omega_{ci,\text{up}}^{-1}\simeq 262\,\text{s}$ in the lab frame, corresponding to a speed below $5\,\text{km/s}$ along x, indicating that the perpendicular shock remains quasi-steady. The time evolution of the total magnetic field strength from $t=10\,\omega_{ci,\text{up}}^{-1}$ to $20\,\omega_{ci,\text{up}}^{-1}$ is depicted in the stack plot (Figure 1f), overlaid with iso-density lines. The overshoots (dark blue) and undershoots (white) are clearly visible across the shock front. Particularly, magnetic and density oscillations within $20\,d_{i,\text{up}}$ near downstream move synchronously with the shock front, suggesting they are stationary structures rather than propagating waves. Further downstream, the slope of the iso-density contours coincides with the plasma bulk velocity, indicating their origin as the structures moving with the plasma flow.
@@
Correlated periodic downstream fluctuations are observed in the plasma density, total magnetic field, and electron thermal pressures (Figure 1a, 1c, and 1d). Ion gyration and the formation of downstream ring distributions (Figure 1i and 1m) are intrinsic consequences of ion dynamics across the shock front. In the downstream region, most ions directly penetrate the shock front and are decelerated by the cross-shock potential. The fraction of reflected ions is small (Figure 1g), limited to a normal length of $2\,d_{i,\text{up}}$ in the ramp. Minimal ion deflection by the magnetic field occurs due to the ramp width being much smaller than the upstream ion convective gyroradius $r_{ci,\text{up}}=V_{i,\text{up}}/\omega_{ci,\text{up}}$. These non-isotropic reflected ions (Figure 1g, 1q-r), after being drift-accelerated along the y-direction, recross the ramp and form a downstream gyrating beam, leading to oscillating ion pressure. The spatial scale of the periodic structure $L$ corresponds to the ion ring distribution in Figure 1i, centered at $V_{ix}=-2V_{Ai,\text{up}},V_{iy}=0$ with a radius of $V_{ring}\simeq 8V_{Ai,\text{up}}$, where $L\sim 2.4\,d_{i,\text{up}}\sim m_{i}V_{ring}/(q_{i}B_{\text{down}})$). Due to their frozen-in behavior with the background magnetic field, the total magnetic field also exhibits coherent periodicity (Figure 1d). With suppressed wave-particle interactions due to the 1D configuration and slow gyrophase mixing, this postshock periodic structure persists nearly to the downstream simulation boundary.
@@
We conducted a companion simulation with $\Delta x=0.5\,d_{e}$ and confirmed that the downstream periodic structures are ion-scale phenomena, independent of electron-scale numerical convergence. However, the inclusion of electron kinetics in FLEKS introduces electrostatic electron pressure modulation by ion and magnetic pressure variations, evident in the correlated spatial oscillation of the $V_{ex}$ extent (Figure 1h) and the modulated pressure components $P_{exx}$, $P_{eyy}$, and $P_{ezz}$ (Figure 1c and 1h).